類数が一定以下の CM 体 (虚 Abel 体のみを除く) の個数の有限性および類数が小さい虚 CM 体 (虚 Abel 体のみを除く) の決定に関する文献

(2012 年 1 月更新版)

[1] R. Brauer : On the zeta functions of algebraic number fields II, Amer. J. Math. 72 (1950), No. 4, 739–746; doi:10.2307/2372290 [PDF (565 KB)], JSTOR; reprinted in Collected Papers, volume III : Finite groups, Lie groups, number theory, polynomials and equations: geometry and biography (P. Fang and W. J. Wong, eds.), MIT Press, Cambridge, Mass.-London, 1980, [58], pp. 489–496; MR0039009 (12, 482g) (reviewed by M. Krasner; reprinted in Reviews in Number Theory, volume 5, R42-6) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0038.17602 (reviewed by Hasse) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften.

[2] K. Uchida : Class numbers of imaginary abelian number fields. I, Tôhoku Math. J. (2) [東北數學雑誌第二輯] 23 (1971), no. 1, 97–104; doi:10.2748/tmj/1178242689, euclid.tmj/1178242689 [PDF (470 KB)]; 23_97.pdf (423 K), Journal@rchive, JST; MR0285509 (44 #2727) (reviewed by J. Browkin; reprinted in Reviews in Number Theory, volume 5, R32-73) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0213.06903 (Autorreferat) [scanned image].

[3] L. J. Goldstein : A generalization of Stark’s theorem, J. Number Theory 3 (1971), no. 3, 323–346; doi:10.1016/0022-314X(71)90008-4 [PDF (917 K)]; MR0289456 (44 #6646) (reviewed by G. Whaples; reprinted in Reviews in Number Theory, volume 5, R14-109) [html, PDF, DVI, PostScript, Bib-TeX]; Zbl 0221.12005 (Autorreferat) [scanned image].

[4] L. J. Goldstein : Relative imaginary quadratic fields of class-number 1 or 2, Trans. Amer. Math. Soc. 165 (1972), 353–364; DOI: 10.1090/S0002-9947-1972-0291124-8 [PDF]; doi:10.2307/1995892 [PDF (266 KB)], JSTOR; MR0291124 (45 #218) (reviewed by T. Kubota; reprinted in Reviews in Number Theory, volume 5A, R14-201) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0245.12007 (reviewed by T. Lepistö) [scanned image].

[5] J. E. Sunley : On the class numbers of totally imaginary quadratic extension of totally real fields, Bull. Amer. Math. Soc. 78 (1972), no. 1, 74–76; DOI: 10.1090/S0002-9904-1972-12859-3 [PDF]; euclid.bams/1183533403 [PDF (282 KB)]; MR0291127 (45 #221) (reviewed by O. F. G. Schilling; reprinted in Reviews in Number Theory, volume 5A, R26-200) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0231.12009 (reviewed by W. Zink) [scanned image].

[6] L. J. Goldstein : Relative imaginary quadratic fields of low class-number, Bull. Amer. Math. Soc. 78 (1972), no. 1, 80–81; DOI: 10.1090/S0002-9904-1972-12864-7 [PDF]; euclid.bams/1183533405 [PDF (167 KB)]; MR0294288 (45 #3357) (reviewed by O. F. G. Schilling; reprinted in Reviews in Number Theory, volume 5A, R26-201) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0232.12001 (reviewed by W. Zink) [scanned image].

[7] J. S. Sunley : Class numbers of totally imaginary quadratic extension of totally real fields, Trans. Amer. Math. Soc. 175 (1973), 209–232; DOI: 10.1090/S0002-9947-1973-0311622-9 [PDF]; doi:10.2307/1996078 [PDF (489 KB)], JSTOR; MR0311622 (47 #184) (reviewed by H. Yokoi; reprinted in Reviews in Number Theory, volume 5A, R26-200E) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0289.12010 (reviewed by T. Lepistö) [scanned image].

[8] J. S. Sunley : Class numbers of totally imaginary fields, Analytic Number Theory (St. Louis Univ., St. Louis, Mo., 1972) (H. G. Diamond, ed.), Proc. Sympos. Pure Math., Vol.

XXIV, Amer. Math. Soc., Providence, R.I., 1973, pp. 303– 304; MR0332722 (48 #11048) (reviewed by Editors; reprinted in Reviews in Number Theory, volume 5A, R26-200F) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0266.12005 (Aus der Arbeit) [scanned image].

[9] K. Uchida : Relative class numbers of normal CMfields, Tôhoku Math. J. (2) [東北數學雑誌第二輯] 25 (1973), no. 3, 347–353; doi:10.2748/tmj/1178241335, euclid.tmj/1178241335 [PDF (421 KB)]; 25_347.pdf (426 K), Journal@rchive, JST; MR0371856 (51 #8073) (reviewed by L. J. Goldstein; reprinted in Reviews in Number Theory, volume 5A, R26-225) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0268.12003 (reviewed by V. Ennola) [scanned image].

[10] H. M. Stark : Some effective cases of the Brauer-Siegel theorem, Invent. Math. 23 (1974), fasc. 2, 135–152; DOI: 10.1007/BF01405166 [u233501484267084.pdf (675.0 KB)];

PPN356556735_0023_log16.pdf (1,15 MB), GDZ; [PDF], Di-

giZeitschriften; MR0342472 (49 #7218) (reviewed by W. Narkiewicz; reprinted in Reviews in Number Theory, volume 5A, R42-220) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0278.12005 (reviewed by A. Yokoyama) [scanned image].

[11] A. M. Odlyzko : Some analytic estimates of class numbers and discriminants, Invent. Math. 29 (1975), fasc. 3, 275–286; DOI: 10.1007/BF01389854 [rq10223725216157.pdf

(497.3 KB)]; PPN356556735_0029_log26.pdf (860,85 KB),

GDZ; [PDF], DigiZeitschriften; MR0376613 (51 #12788) (reviewed by L. J. Goldstein; reprinted in Reviews in Number Theory, volume 5A, R28-207) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0299.12010 (Autorreferat) [scanned image], 306.12005 [scanned image].

[12] H. M. Stark : The analytic theory of algebraic integers, Bull. Amer. Math. Soc. 81 (1975), no. 6, 961– 972; DOI: 10.1090/S0002-9904-1975-13873-0 [PDF]; euclid.bams/1183537391 [PDF (1009 KB)]; MR0444611 (56 #2961) (reviewed by E. Grosswald; reprinted in Reviews in Number Theory, volume 5A, R42-253) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0329.12010 (reviewed by V. Şt. Udrescu) [scanned image].

[13] J. B. Friedlander : On the class numbers of certain quadratic extensions, Acta Arith. 28 (1975), fasc. 4, 391– 393; aa2846.pdf, icm; MR0389853 (52 #10683) (reviewed by W. Narkiewicz) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0329.12011 (reviewed by T. Metsänkylä) [scanned image].

[14] J. Hoffstein : Some analytic bounds for zeta functions and class numbers, Invent. Math. 55 (1979), fasc. 1, 37–47; DOI: 10.1007/BF02139701 [n853256g51175310.pdf (418.0 KB)];

PPN356556735_0055_log8.pdf (732,92 KB), GDZ; [PDF],

DigiZeitschriften; MR0553994 (80k:12019) (reviewed by H. Iwaniec; reprinted in Reviews in Number Theory, volume 5A, R42-278) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0474.12009 (reviewed by S. Uchiyama) [scanned image].

[15] J. Hoffstein and N. Jochnowitz : On Artin’s conjecture and the class number of certain CM-fields, I, Duke Math. J. 59 (1989), no. 2, 553–563; DOI: 10.1215/S0012-7094-89-05925-5, euclid.dmj/1077308015 [PDF]; MR1016903 (90h:11104) (reviewed by D. Goldfeld; reprinted in Reviews in Number Theory, volume 5B, R29-771) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0711.11042 (reviewed by S. Kanemitsu) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[16] J. Hoffstein and N. Jochnowitz : On Artin’s conjecture and the class number of certain CM-fields, II, Duke Math. J. 59 (1989), no. 2, 565–584; DOI: 10.1215/S0012-7094-89-05926-7, euclid.dmj/1077308016 [PDF]; MR1016903 (90h:11104) (reviewed by D. Goldfeld; reprinted in Reviews in Number Theory, volume 5B, R29-771) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0711.11042 (reviewed by S. Kanemitsu) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[17] S. Louboutin : On the class number one problem for nonnormal quartic CM-fields, Tôhoku Math. J. (2) [東北數學雑誌第二輯] 46 (1994), no. 1, 1–12; doi:10.2748/tmj/1178225798, euclid.tmj/1178225798 [PDF (947 KB)]; 46_1.pdf (461 K), Journal@rchive, JST; MR1256724 (94m:11130) (reviewed by K. Yamamura; reprinted in Reviews in Number Theory, volume 5B, R29-894) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0796.11050 (reviewed by S. Louboutin) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[18] S. Louboutin and R. Okazaki : Determination of all nonnormal quartic CM-fields and of all non-abelian normal octic CM-fields with class number one, Acta Arith. LXVII (1994), fasc. 1, 47–62; aa6714.pdf, icm; MR1292520 (95g:11107) (reviewed by K. Yamamura; reprinted in Reviews in Number Theory, volume 5B, R21-635) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0809.11069 (reviewed by S. Louboutin) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[19] S. Louboutin and R. Okazaki : Determination of all quaternion octic CM-fields with ideal class groups of exponents 2. Abridged version, Proc. Japan Acad. Ser. A Math. Sci. [日本学士院紀要] 70 (1994), No. 2, 56–61; doi:10.3792/pjaa.70.56, euclid.pja/1195511137 [PDF (616 KB)]; MR1272671 (95d:11152) (reviewed by H. Yokoi; reprinted in Reviews in Number Theory, volume 5B, R29-904) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0809.11070 (summary) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[20] S. Louboutin : The CM-fields with class number one which are Hilbert class fields of quadratic number fields, Proc. Japan Acad. Ser. A Math. Sci. [日本学士院紀要] 72 (1996), No. 1, 16–19; doi:10.3792/pjaa.72.16, euclid.pja/1195510481 [PDF (513 KB)]; MR1382779 (97b:11131) (reviewed by M. Olivier) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0854.11059 (reviewed by R. Mollin) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[21] S. Louboutin : Corps quadratiques à corps de classes de Hilbert principaux et à multiplication complexe, Acta Arith. LXXIV (1996), fasc. 2, 121–140; aa7423.pdf, icm; MR1373703 (96m:11098) (reviewed by F. Diaz y Diaz; reprinted in Reviews in Number Theory, volume 5B, R29-950) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0854.11058 (reviewed by S. Louboutin) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[22] S. Louboutin : Determination of all quaternion octic CM-fields with class number 2, J. London Math. Soc.

(2) 54 (1996), Part 2, 227–238; doi:10.1112/jlms/54.2.227 [227.full.pdf]; MR1405052 (97g:11122) (reviewed by K. Yamamura) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0861.11064 (reviewed by S. Louboutin) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[23] S. Louboutin, R. Okazaki, and M. Olivier : The class number one problem for some non-abelian normal CM-fields, Trans. Amer. Math. Soc. 349 (1997), no. 9, 3657–3678; DOI: 10.1090/S0002-9947-97-01768-6 [PDF]; doi:10.2307/2155534 [PDF (1938 KB)], JSTOR; MR1390044 (97k:11149) (reviewed by B. de Smit) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0893.11045 (reviewed by E. Benjamin) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[24] S. Louboutin : The class number one problem for the nonabelian normal CM-fields of degree 16, Acta Arith. LXXXII (1997), fasc. 2, 173–196; aa8225.pdf, icm; MR1477509 (98j:11097) (reviewed by K. Yamamura) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0881.11079 (reviewed by S. Louboutin) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[25] S. Louboutin and R. Okazaki : The class number one problem for some non-abelian normal CM-fields of 2-power degrees, Proc. London Math. Soc. (3) 76 (1998), Part 3, 523–548; doi:10.1112/S0024611598000318

                                                            [S002461159800031Xa.pdf (370 Kb), CJO; 523.full.pdf,

Oxford Journals]; MR1616805 (99c:11138) (reviewed by M. Olivier) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0891.11054 (reviewed by S. Louboutin) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[26] S. Louboutin, Y.-H. Park and Y. Lefeuvre : Construction of the real dihedral number fields of degree 2p. Applications, Acta Arith. LXXXIX (1999), fasc. 3, 201–215; aa8931.pdf, icm; MR1691851 (2000g:11101) (reviewed by M. Olivier) [html, PDF, DVI, PostScript, BibTeX].

[27] S. Louboutin : The class number one problem for the dihedral and dicyclic CM-fields, Colloq. Math. 80 (1999), No. 2, 259–265; cm8029.pdf, icm; MR1703822 (2000e:11140) (reviewed by M. Olivier) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1036.11056 (reviewed by F. Lemmermeyer) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[28] H.-S. Yang and S.-H. Kwon : The non-normal quartic CMfields and the octic dihedral CM-fields with relative class number two, J. Number Theory 79 (1999), no. 2, 175– 193; doi:10.1006/jnth.1999.2421 [PDF (161 K)]; MR1728146 (2000h:11117) (reviewed by F. Lemmermeyer) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0976.11051 (reviewed by H. Yokoi) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[29] F. Lemmermeyer, S. Louboutin, et R. Okazaki : The class number one problem for some nonabelian normal CM-fields of degree 24, J. Théor. Nombres Bordeaux 11 (1999), Fasc. 2, 387–406;

                                                            doi:  10.5802/jtnb.257 [JTNB_1999__11_2_387_0.pdf];
                                                            JTNB_1999__11_2_387_0.djvu,  JTNB_1999__11_2_

387_0.pdf, Numdam; MR1745886 (2001j:11104) (reviewed by K. Yamamura) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1010.11063 (summary) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[30] S. Louboutin and R. Okazaki : Determination of all quaternion CM-fields with ideal class groups of exponents 2, Osaka J. Math. 36 (1999), no. 2, 229–257; euclid.ojm/1200788566 [PDF (2260 KB)]; ojm36_02_01.pdf, URL: http://ir.library.osaka-u.ac.jp/meta-bin/mt-pdetail.cgi ?smode=1&edm=0&tlang=1&cd=00028474; MR1736479 (2001c:11120) (reviewed by M. Olivier) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0952.11025 (reviewed by F. Lemmermeyer) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[31] R. Okazaki : Non-normal class number one problem and the least prime power-residue, Number theory and its applications (Kyoto, 1997) (S. Kanemitsu and K. Győry, eds.), Dev. in Math., vol. 2, Kluwer Acad. Publ., Dordrecht, 1999, pp. 263–272; MR1738823 (2000k:11125) (reviewed by S. Louboutin) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0972.11106 (reviewed by F. Lemmermeyer) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[32] 山村健 [Yammura, K.] : A survey of small class number problems for CM-fields and related problems, 解析数論と数論諸分野の交流 [Number Theory from the Stand Point of Analytic Number Theroy] (京都大学数理解析研究所, 1998) (中井喜信編), 数理解析研究所講究録 [RIMS Kokyuroku] 1091, 1999, 144–156; 1091-17.pdf; 1091-17.pdf (1.15 MB), URI: http://hdl.handle.net/2433/62900; MR1744902.

[33] 山村健 : A survey of small class number problems for CMfields and related problems, 内田興二先生還暦記念研究集会 (東北大学, 1999), 仙台数論小研究集会 1999 報告集 (内田興二編), 東北大学, 2000, pp. 13–20; yamamura.dvi.

[34] Y. Lefeuvre : Corps diédraux à multiplication comlexe principaux, Ann. Inst. Fourier (Grenoble) 50 (2000),

                                                            Fasc. 1, 67–103; AIF_2000__50_1_67_0.pdf, Cedram;
                                                            AIF_2000__50_1_67_0.djvu,  AIF_2000__50_1_67_0.pdf,

Numdam; MR1762338 (2001g:11166) (reviewed by M. Olivier) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0952.11024 (reviewed by V. Ennola) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[35] S. Louboutin and Y.-H. Park : The class number problem for dicyclic CM-fields, Publ. Math. Debrecen 57 (2000), no. 3–4, 283–295; [PDF]; MR1798714 (2001m:11196) (reviewed by R. Okazaki) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0963.11065 (reviewed by C. Greither) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[36] S. Louboutin : Explicit bounds for residues of Dedekind zeta functions and values of L-functions ar s = 1, and relative class numbers, J. Number Theory 85 (2000), no. 2, 263– 282; doi:10.1006/jnth.2000.2545 [PDF (168 K)]; MR1802716 (2002i:11111) (abstract) [html, PDF, DVI, PostScript, Bib-TeX]; Zbl 0967.11049 (reviewed by F. Lemmermeyer) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[37] Y.-H. Park : On the class number problem for normal CM-fields, Trends in Math. 3 (2000), 93–101; 974361.dvi,

974361.pdf.

[38] Y. Lefeuvre and S. Louboutin : The class number one problem for dihedral CM-fields, Algebraic number theory and Diophantine analysis (Graz, 1998) (F. Halter-Koch and R. F. Tichy, eds.), de Gruyter, Berlin, 2000, pp. 249–275; [eBook]; MR1770466 (2001k:11224) (reviewed by M. Olivier) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0958.11071 (reviewed by F. Lemmermeyer) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[39] S. Louboutin : Explicit bounds for residues of Dedekind zeta functions and values of L-functions ar s = 1, and relative class numbers (abridged version), 解析的整数論とその周辺 [Analytic Number Theory and Related Topics] (京都大学数理解析研究所, 1999) (松本耕二編), 数理解析研究所講究録 [RIMS Kokyuroku] 1160, 2000, 210–219; 1160-28.pdf; 1160-28.pdf (834.42 kB), URI: http://hdl.handle.net/2433/64226; MR1799164 (not reviewed); Zbl 0969.11547 (not reviewed).

[40] S. Louboutin : Computation of L(0, χ) and of relative class numbers of CM-fields, Nagoya Math. J. 161 (2001), 171– 191; euclid.nmj/1114631557 [PDF (195 KB)]; MR1820217 (2002e:11152) (reviewed by R. Okazaki) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0985.11054 (reviewed by O. Ramaré) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[41] S. Louboutin : Explicit upper bounds for residues of Dedekind zeta functions and values of L-functions ar s = 1, and explicit lower bounds for relative class numbers of CM-fields, Canad. J. Math. 53 (2001), no. 6, 1194– 1222; DOI: 10.4153/CJM-2001-045-5 [louboutin1046.pdf [256 KB]]; MR1863848 (2003d:11167) (abstract) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0998.11066 (reviewed by F. Lemmermeyer) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[42] V. K. Murty : Class numbers of CM-fields with solvable normal closure, Compositio Math. 127 (2001), no. 3, 273–288; doi:10.1023/A:1017589432526 [n683276rm7542jp1pdf.pdf (141.4 KB), Springer; S0010437X01000458a.pdf (144 K), CJO]; MR1845038 (2003a:11147) (reviewed by T. Funakura) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1116.11091 (reviewed by O. Ninnemann) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[43] K.-Y. Chang and S.-H. Kwon : The non-abelian normal CM-fields of degree 36 with class number one, Acta Arith. 101 (2002), fasc. 1, 53–61; doi:10.4064/aa101-1-4 [pdf_aa101-1-04.pdf (151.38 Kb)]; MR1879845 (2003e:11119) (reviewed by O. Ramare) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0998.11059 (reviewed by O. Ramaré) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[44] Y.-H. Park : The class number one problem for the non-abelian normal CM-fields of degree 24 and 40, Acta Arith. 101 (2002), fasc. 1, 63–80; doi:10.4064/aa101-1-5 [pdf_aa101-1-05.pdf (227.71 Kb)]; MR1879844 (2002k:11200) (reviewed by K. Yamamura) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0997.11091 (reviewed by O. Ramaré) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[45] G. Boutteaux and S. R. Louboutin : The class number one problem for some non-normal sextic CM-fields, Analytic number theory (Beijing/Kyoto, 1999) (C. Jia and K. Matsumoto, eds.), Dev. Math., vol. 6, Kluwer Acad. Publ., Dordrecht, 2002, pp. 27–37; MR1901973 (2003c:11146) (reviewed by T. Metsänkylä) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1023.11057 (reviewed by F. Lemmermeyer) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[46] G. Boutteaux and S. R. Louboutin : The class number one problem for some non-normal sextic CM-fields. Part 2, Acta Math. Info. Univ. Ostraviensis 10 (2002), no.

                                                            1, 3–23; ActaOstrav_10-2002-1_1.pdf (3.2 MB), DML-

CZ; MR1943019 (2003i:11162) (reviewed by T. Metsänkylä) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1042.11070 (reviewed by F. Lemmermeyer) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[47] S. Bessassi : Bounds for the degree of CM-fields of class number one, Acta Arith. 106 (2003), fasc. 3, 213–245; doi:10.4064/aa106-3-2 [pdf_aa106-3-02.pdf (248.58 Kb)]; MR1957106 (2003m:11183) (reviewed by K. Yamamura) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1146.11329 (reviewed by K. Yamamura) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[48] K.-Y. Chang and S.-H. Kwon : The class number one problem for some non-abelian normal CM-fields of degree 48, Math. Comp. 72 (2003), no. 242, 1003–1017; DOI: 10.1090/S0025-5718-02-01443-6 [PDF]; doi:10.2307/4099945 [PDF (475 KB)], JSTOR; MR1954981 (2003m:11185) (reviewed by S. Louboutin) [html, PDF, DVI, PostScript, Bib-TeX]; Zbl 1071.11062 (reviewed by K. Yamamura) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[49] R. Okazaki : Class number one problem in non-normal sextic CM-fields, Proceedings of the 2003 Nagoya Conference “Yokoi-Chowla Conjecture and Related Problems” (Nagoya, 2003) (S.-I. Katayama, C. Levesque and T. Nakahara, eds.), Saga Univ., Saga, 2004, pp. 101–104; MR2109027 (2005k:11223) (reviewed by A. Mouhib) [html, PDF, DVI, PostScript, BibTeX].

[50] G.-N. Lee and S.-H. Kwon : CM-fields with relative class number one, Math. Comp. 75 (2006), no. 254, 997–1013; DOI: 10.1090/S0025-5718-05-01811-9 [PDF]; doi:10.2307/4100323 [PDF], JSTOR; MR2197004 (2006k:11216) (reviewed by J.- R. Belliard) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1143.11353 (summary) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[51] S.-M. Park and S.-H. Kwon : Class number one problem for normal CM-fields, J. Number Theory 125 (2007), no. 1, 59– 84; doi:10.1016/j.jnt.2006.10.012 [PDF (360 K)]; MR2333119 (2008c:11149) (reviewed by S. Louboutin) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1128.11053 (reviewed by C. Parry) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[52] H.-S. Yang, Y.-H. Park, and S.-H. Kwon : Class number one problem for normal CM-fields of degree 32, Trans. Amer. Math. Soc. 359 (2007), no. 10, 5057–5089; DOI: 10.1090/S0002-9947-07-04219-5 [PDF]; MR2320660 (2008f:11129) (reviewed by T. Metsänkylä) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1162.11392 (summary) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[53] G.-N. Lee and S.-H. Kwon : The zeros of Dedekind zeta functions and class numbers of CM-fields, Math. Comp. 77 (2008), no. 264, 2437–2445; DOI: 10.1090/S0025-5718-08-02093-0 [PDF]; MR2429892 (2009g:11150) (reviewed by J. Kaczorowski) [html, PDF, DVI, PostScript, BibTeX]; Zbl pre05813029 (reviewed by T. Metsänkylä) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[54] S.-H. Kwon, S. Louboutin and S.-M. Park : Nonabelian normal CM-fields of degree 2pq, J. Aust. Math. Soc. 87 (2009), Part 1, 129–144; doi:10.1017/S1446788709000081 [PDF (203 KB)]; 48188499.pdf (211 KB), Academic Search Complete, EBSCO; MR2538641 (2010g:11188) (reviewed by R. Quême) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1196.11151 (reviewed by F. Lemmermeyer) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[55] S.-H. Kwon : Class numbers of CM-fields, 第 8 回北陸数論研究集会報告集 ( 「ガロア理論とその周辺」とのジョイント開催) [Proceedings of the 8th Workshop on Number Theory] (金沢大学サテライトプラザ, 2009) (平林幹人, 野村明人, 山下浩, 木村巌, 菅野孝史編), 2010, pp. 97–110.

[56] 山村健 : 類数が一定以下の CM 体 (虚 Abel 体のみを除く) の個数の有限性および類数が小さい虚 CM 体 (虚 Abel 体のみを除く) の決定に関する文献 (2010 年 4 月更新版), 第 8 回北陸数論研究集会報告集 ( 「ガロア理論とその周辺」とのジョイント開催) [Proceedings of the 8th Workshop on Number Theory] (金沢大学サテライトプラザ, 2009) (平林幹人, 野村明人, 山下浩, 木村巌, 菅野孝史編), 2010, pp. 238–245.

[57] S. Louboutin and Y.-H. Park : The class number problem for non-abelian normal CM-fields of degree 42, preprint.

[58] Y.-H. Park, and S.-H. Kwon : Class number one problem for normal CM-fields of degree 32. II, preprint, 2001.

[59] R. Okazaki : On class number one problem in non-normal sextic CM-fields with imaginary quadratic subfields, preprint;

NNcubicIQ.pdf, NNcubicIQ.ps, NNcubicIQ,dvi.

[60] R. Okazaki : On class number one problem in non-normal sextic CM-fields with imaginary quadratic subfields, preprint;

NNcubicIQ.pdf, NNcubicIQ.ps, NNcubicIQ.dvi.